Arginine is an amino acid which has the following pKa values: pK1: 2.17, pK2: = 9.04, pKR = 12.48. Calculate the net charge of arginine at the pH value of 4.
1
Expert's answer
2021-08-30T02:26:31-0400

pH = (pKa1 + pKa2)/2.

Therefore, pH=pKa=(pK1+pK2+pKR)/3

=(2.17+9.04+12.48)/3

pH=7.9

According to the below referencing material, the Arginine's net charge is +1




The above picture is from https://www.mtholyoke.edu/courses/sdecatur/chem333/problems/ps1key.html